OpenStudy (anonymous):

using factoring and the zero product property to find the zeros of each quadratic function. f(x)=4x^2+4x-24

OpenStudy (campbell_st):

factor out 4 \[f(x) = 4(x^2 +x - 6)\] now factor the quadratic.... hope it helps

OpenStudy (anonymous):

i got that far, but like i dont really know what to do afterwards.

OpenStudy (campbell_st):

ok... so what are the factors of -6 that add to 1 the larger factor is positive... and the smaller negative so f(x) =4(x + factor1)(x + factor 2)

OpenStudy (anonymous):

so then, f(x)=4(x+3)(x-2)?

OpenStudy (campbell_st):

yep... now to find the zeros solve x + 3 = 0 x -2 = 0

OpenStudy (anonymous):

oooh okay thnak yooou ^_^